OpenStudy (anonymous):

9x-3y≤-12 in slope intercept form

OpenStudy (anonymous):

slope intercept form is just y=mx+b or in this case, \[y \le mx+b\] does that help?

OpenStudy (anonymous):

9x + 3y >= -12 3y >= -9x - 12 y >= -(9/3)x - 12/3 y >= -3x - 4 slope is -3, intercept is -4 Hope that helps!